OpenStudy (anonymous):

(Algebra 1 Help!!) Mary wants to hang a mirror in her room. The mirror and frame must have an area of 7 square feet. The mirror is 2 feet wide and 3 feet long. Which quadratic equation can be used to determine the thickness of the frame, x?

OpenStudy (anonymous):

Here are my option: x2 + 14x - 2 = 0 2x2 + 10x - 7 = 0 3x2 + 12x - 7 = 0 4x2 + 10x - 1 = 0

OpenStudy (anonymous):

If x is the thickness of the frame, then 2x + 2 is the width and 2x + 3 is the length of the mirror with frame. Area of the frame now is 2x + 2 times 2x + 3, which should equal 7 So the equation becomes (2x + 2)(2x + 3) = 7 Simplifying it you get 4x^2 + 10x - 1 = 0

OpenStudy (anonymous):

@navk Thank you, friend.